# Here Be Dragons: Bimodal posteriors arise from numerical integration error in longitudinal models

### v2

@@ 2.1 Numerical integration theory
We have described an example case where $\beta_{1}=1$, and $\beta_{0}=10$. The two real solutions to Equation (2.19) for those values are $\hat{\beta}_{1}=1.0008$ and $4.9112$, which is shown in Figure 2 and aligns with the behaviour in Figure 1. The small difference from $\beta_{1}=1$ in the better estimate is to be expected given Corollary 2.4.
[…]
To conclude the theory section, the posterior distribution for $\beta_{1}$ and $\beta_{0}$ conditioned on the RK4 step will have modes at the parameter values that produce $\hat{Y}(t_{j})$ that match the analytic solution. The solutions to Equation (2.19) gives posterior modes that cannot be distinguished by estimation based on a longitudinal model using the RK4 algorithm, which is precisely what non-identifiability means in this context.
Equation 2.19 is calculable directly because we have an analytic solution, explicit numerical method, and a straight-forward function. Attempts to do the same with the non-linear Canham function (Canham et al.,, 2004) for example would be impossible as there is no analytic solution for comparison. While the substitution with the analytic solution is not essential to testing a step where a better numerical method can give $Y(t_{j}+h)$, it would be intractable for Canham anyway due to the non-linear DE. In general it may not be possible to do an explicit formula of that form. Simulation is much easier and the user can test different methods and estimation tools as well.
@@ 2.2 Aims
We have three questions about the pathological behaviour:
  • Do different integration methods have persistent bimodality? We have demonstrated mathematical bimodality for RK4 with a step size of 0.5 so we test whether that shows up empirically as well. We also do empirical tests for RK4 with step sizes 0.25 and 0.125, RK45 with adaptive step size, and an analytic solution.
@@
  • Is there a model that provides accurate unimodal estimates? We do simulation studies for the affine and Canham models to test whether the chosen work-arounds – an analytic implementation and RK45 with adaptive step size respectively – have reliable results.
@@ 2.3 Methods
To address the questions we simulated data so we know the true values. We produce ‘true’ data from the analytic solution Equation (2.3) with parameters $\beta_{0}=10$, $\beta_{1}=1$ and initial size $Y(0)=1$. The simulated survey structure is 10 size values $Y(t_{j})$ with an interval of 1 unit of time between them, so $t_{j+1}-t_{j}=1$. From the analytic sizes over time we get ‘observations’ with error, rounded to precision 0.1, of the form
@@
[…]
@@
To avoid convergence problems we initialise the finite mixture algorithm with the sample means of the two initial clusters. The fitted mixture of normal distributions is then used to analyse the clustering behaviour, as we get cluster mean estimate vectors, covariance matrices, and cluster occurrence probability estimates. We also use scatter plots to look at the cluster marginal posterior parameter distributions. In the unimodal case we look at the mean of the estimated posterior parameters to check for evidence of large, and a 95% posterior credible interval (Gelman et al.,, 2021) built from taking the central 95% quantile interval of the posterior estimates to get an idea of spread.
@@
2.3.1 Checking Canham
We use simulation to check whether the Canham model is robust to posterior clustering due to numerical error with the RK45 method and adaptive step size. The differential equation defined in Canham et al., (2004) is
$f(Y(t),g_{max},y_{max},k)=g_{max}\exp\Bigg{(}-\frac{1}{2}\bigg{(}\frac{\ln(Y(t)/y_{max})}{k}\bigg{)}^{2}\Bigg{)},$ (2.20)
which is extremely non-linear and has no analytic solution, necessitating numerical methods within the longitudinal model. Equation (2.20) has advantages for numerical estimation over the affine ODE: it is strictly positive which helps avoid the extreme oscillation we see in Figure 1 from negative growth, and it is bounded above by $g_{max}$ which limits how big positive growth increments can get. The affine function is unbounded above and below, and because a negative growth increment can lead to negative $Y(t)$ values even if the data is strictly positive (Canham is undefined for $Y(t)\leq 0$ due to the log function), it is possible for the numerical method to encounter extreme positive growth increments as well as negative ones.
Table 1: Posterior finite mixture model estimates for MCMC testing different numerical methods for the affine ODE. True parameter values are $\beta_{0}=10$ and $\beta_{1}=1$. The sample standard deviations of each cluster for each parameter are given in parentheses. Note that for each choice of step size, RK4 converged to one of two very different solutions.
Cluster 1 Cluster 2 Cluster 2
Numerical method Step size $\mu_{\beta_{0},1}\ (sd(\hat{\beta}_{0},1))$ $\mu_{\beta_{1},1}\ (sd(\hat{\beta}_{1},1))$ $\mu_{\beta_{0},2}\ (sd(\hat{\beta}_{0},2))$ $\mu_{\beta_{1},2}\ (sd(\hat{\beta}_{1},2))$ prob.
@@ 2.3.1 Checking Canham
There is no analytic solution to Equation (2.20) which would give access to ‘true’ sizes over time $Y(t)$. Instead we used the deSolve RK4 method with step size 0.0001 to project forward 50 units of time from an initial size of $Y(0)=1$. The true parameters $g_{max}=0.8,\ y_{max}=8,\ k=1$ were chosen as plausible values based on what was observed in O’Brien et al., (2024). To test whether the step size was small enough to give good numerical estimates, we did the same forward projection with step size 0.001 and compared the two, which gave a maximum difference of $3.02\times 10^{-13}$. We fit 10,000 chains to ‘observed’ sizes: from the smaller step size, we took 10 points, 5 years apart as the true values, then for each chain we add independent error from $\mathcal{N}(0,0.1)$, and rounded to precision 0.1. We added error independently for each chain as We are interested in whether the Canham model has evidence of bias as well as multi-modality.
We use visualisations of the distribution of posterior means. Scatter plots and histograms are used to examine clustering. We use the mean of the estimates and central 95% posterior credible interval to examine spread and bias. The posterior is unimodal and attempts to fit a mixture distribution with two clusters failed as a result, so we do not use one.
3 Affine DE results
3.1 Numerical methods have persistent bimodality
All the step sizes tested with the RK4 algorithm had a second cluster of erroneous parameter estimates. As the step size shrinks, the erroneous cluster moves further away from the true values, and the probability of erroneous convergence decreases as shown in Table 2.3.1. Figure 3 shows the cluster distance in Panel (b), where all 30,000 estimates are plotted together yet they appear as if single points due to the tiny within-cluster variance, as in Table 2.3.1. There is a stack of points at (9.95, 0.991) which is the first cluster of estimates for all step sizes, and each step size has a second cluster spread further from that point as the step size decreases. The clusters all lie along a line preserving the asymptotic size at $\hat{\beta}_{0}/\hat{\beta}_{1}\approx 10$, which is also visible in Figure 3(c) and (d). Panel (c) shows Equation (2.1) with parameters fit to the mean of each erroneous cluster compared to the true values, and Panel (d) shows the analytic solution Equation (2.3) with the same. The erroneous cluster has much faster analytic solution convergence to the asymptotic size, but as we saw in Figure 1, that leads to bad numerical behaviour.
Table 2: Posterior finite mixture model estimates and prior densities for MCMC with different priors. The prior density is the density of the prior distribution at the cluster centre and indicates relative probability before model fitting, while posterior probability is based on the number of points in each cluster. True parameter values are $\beta_{0}=10$ and $\beta_{1}=1$. Numerical integration used the RK4 method with a step size of 0.5. The $\beta_{c}$ prior mean includes the $\bar{y}$ shift, and the $\beta_{1}$ mean has a log transformation. Note that results, where the posterior mean was one of two very different solutions (each with high probability), are robust to prior choice, with the second solution only disappearing when an extremely tight prior was chosen around the first solution.
Posterior locations
Cluster 1 Cluster 2 Prior density Posterior prob.
Priors $\mu_{\beta_{0},1}$ $\mu_{\beta_{1},1}$ $\mu_{\beta_{0},2}$ $\mu_{\beta_{1},2}$ Cluster 1 Cluster 2 Cluster 1 Cluster 2
Good estimate centre $\beta_{c}\sim\mathcal{N}(1.43,2)$ 9.957 0.9918 49.27 4.914 0.03979 0.0005852 0.6871 0.3129
$\beta_{1}\sim\log\mathcal{N}(-0.00904,2)$
Bad estimate centre $\beta_{c}\sim\mathcal{N}(7.09,2)$ 9.969 0.9928 49.37 4.917 0.0005286 0.03979 0.6822 0.3178
$\beta_{1}\sim\log\mathcal{N}(1.59,2)$
Midpoint $\beta_{c}\sim\mathcal{N}(4.26,2)$ 9.963 0.9923 49.32 4.915 0.0126 0.01456 0.6805 0.3195
$\beta_{1}\sim\log\mathcal{N}(1.09,2)$
Large SD $\beta_{c}\sim\mathcal{N}(1,10)$ 9.957 0.9917 49.370 4.9170 1.590e-3 1.306e-3 0.6842 0.3158
$\beta_{1}\sim\log\mathcal{N}(0,10)$
Small SD $\beta_{c}\sim\mathcal{N}(1,0.1)$ 9.765 0.976 41.65 4.581 0.0153 1.085e-83 0.7437 0.2563
$\beta_{1}\sim\log\mathcal{N}(0,0.1)$
Silly small SD $\beta_{c}\sim\mathcal{N}(1,0.001)$ 9.599 0.9999 9.599 0.9999 149400 149400 0.4072 0.5928
$\beta_{1}\sim\log\mathcal{N}(0,0.001)$
Figure 5: Estimated $\hat{Y}(t)$ across three step sizes compared to the true model for simulated data, and numerical error. (a), (c), and (e) show that the estimated sizes over time for the erroneous cluster align closely with the analytic solution for the true $\beta$s across the step sizes. The slight difference in asymptotic size for step size 0.125 is likely due to the prior constraining $\hat{\beta}_{0}$. (b), (d), and (f) show the numerical error for each erroneous cluster, where the initial step has the largest error and then there is convergence to the same asymptotic size.
We tested a better numerical method as well – RK45 with adaptive step size – and found that under some prior specifications bimodality still occurred. Figure 6 shows unimodal posteriors for both the analytic and RK45 posteriors when $\beta_{c}$ has a normal prior, while Figure 7 shows the bimodal posterior that arose when we was testing a log-normal distribution for $\beta_{c}$ with RK45, a problem that the analytic solution does not have. Log-normal priors are common so we include this example as a warning. Table 2.3.1 gives the mixture-model analysis for RK45. We do not consider the two clusters for $\beta_{c}\sim\log\mathcal{N}$ to be meaningfully different as they are so close together.
@@ 3.1 Numerical methods have persistent bimodality
3.2 Changes to estimation methods do not address the underlying issue
Table 3.1 shows that all alternate priors except for one with an unreasonably small standard deviation picked up an erroneous second mode. Changing the prior central location had minimal impact, as can be seen in the first three prior configurations in Table 3.1. The ‘good’ modes had very similar locations, and the probability of converging to the second (erroneous) mode was not substantially different across the fits. The behaviour of the different centred priors is very similar to the RK4 with 0.5 step size and default prior as seen in Table 2.3.1. That the posterior probability did not change even when the prior was centred at one or the other cluster indicates that changing the prior location does not help with constraining erroneous convergence.
A larger prior standard deviation of 10 produced results that were not meaningfully different from the default of 2. Reducing the prior standard deviation to 0.1 changed where the second mode was, but did not meaningfully constrain its occurrence with a reduction from 32% to 26%. The even more unreasonably small prior standard deviation (0.001) was able to eliminate the second mode as the distance between the two means from the mixture model was vanishingly small. The purpose of including an unreasonable standard deviation prior was to demonstrate that it is mathematically possible to constrain where the posterior goes, but to do so requires something that would not be used in practice.
The deterministic algorithm L-BFGS performed badly overall, as seen in Figure 8. The figure represents 10,000 fits. There were problems where the randomly selected starting point failed to converge for a substantial region of the parameter space. Where the L-BFGS algorithm did converge, it had bimodal estimates.
Figure 8: Estimation with L-BFGS optimisation algorithm. The deterministic method with the RK4 algorithm failed to converge in 5.8% of runs that randomly started in a pathological region as indicated by the blue crosses.
3.3 Our chosen methods produce accurate, unimodal estimates
The simulation tests for the affine model with an analytic solution produced a unimodal distribution that was centred within 0.3% of the true values as shown in Table 3 which is a negligible empirical bias. Using the analytic solution means we do not introduce the non-identifiability that arises from numerical error so this result is in line with our theory.
4 Canham results
The overall behaviour of the posterior estimation for Canham is shown in Figure 9. We have a unimodal posterior distribution with no evidence of additional clusters in any dimension. Parameter estimation information is given in Table 3. We can be reasonably confident that the Canham function is not subject to the same pathologies as Equation (2.1) given the RK45 algorithm with adaptive step size used in the longitudinal model.
The associations between $g_{max}$ and $k$, $y_{max}$ and $k$ in Figure 9(g) and (h) have an interpretation that works with the shape of the function in Panel (b): as $k$ increases, if $g_{max}$ decreases you will get similar total growth but more spread out over the observation period; and for $k$ and $y_{max}$ a negative relationship means that the hump will be more pronounced if $y_{max}$ is large. We suspect that there is a connection between the sampling design (10 observations, 5 units of time apart) and how well the peak can be observed. Investigating that is outside the scope of this paper.
Table 3: The affine model with an analytic solution and the Canham model with RK45 and an adaptive step size show minimal evidence of bias.
@@ 4 Canham results
Affine $\beta_{0}$ 10 10.02 (9.484, 10.57)9.48410.57(9.484,\ 10.57)( 9.484 , 10.57 )
$\beta_{1}$ 1 1.002 (0.9445, 1.062)0.94451.062(0.9445,\ 1.062)( 0.9445 , 1.062 )
Canham $g_{max}$ 0.8 0.799 (0.776, 0.823)0.7760.823(0.776,\ 0.823)( 0.776 , 0.823 )
$y_{max}$ 8 7.979 (7.642, 8.301)7.6428.301(7.642,\ 8.301)( 7.642 , 8.301 )
$k$ 1 1.009 (0.952 1.073)0.9521.073(0.952\ 1.073)( 0.952 1.073 )
Figure 9: (a) shows an example of how the observations compare to the high precision numerical solution given $\mathcal{N}(0,0.1)$ error and measurement precision of 0.1. (b) compares the estimated and true growth functions which demonstrate strong agreement. (c), (d), and (e) histograms of the posterior parameter estimates demonstrate the unimodal posterior distribution and show no evidence of bi- or multi-modality. (c), (d), and (e) pair scatter plots of the posterior parameter estimates show no evidence of multi-modality, and moderate associations between $g_{max}$ and $k$, $y_{max}$ and $k$. The scatter plots and histograms are aligned on the horizontal parameter axis for ease of comparison.
@@ 5 Discussion
This paper provides theoretical and empirical evidence of a multi-modal posterior arising from numerical error in a longitudinal model for an affine first order ODE. One posterior mode aligns with good estimates of the true parameter values for simulated data, and at that mode the numerical method is reasonably well behaved. The other mode has extremely bad estimates of the true parameters, and badly behaved numerics as well. The existence of non-identifiability for RK4 in the case of an affine ODE demonstrates that care must be taken when using numerical methods for longitudinal models.
We show that the affine model can use an analytic solution to avoid the numerical pathology, and also that in the case of the Canham model where there is no analytic solution to use, the choice of RK45 with adaptive step size is sufficient as we have no evidence of multi-modal posteriors. The simulation testing procedure is more generally applicable and should form part of longitudinal model implementation if there is evidence of bad posterior behaviour such as divergent chains. It is also possible to test estimates from real world data after the fact by using a more accurate numerical method than the one implemented in the model.
Using numerical theory, we proved that in general, the RK4 algorithm in a longitudinal model for affine first-order ODEs with negative slope will produce multi-modal results, and demonstrate with a specific ODE. In the empirical results, we see two clusters of parameter estimates that produce statistically indistinguishable $\hat{Y}(t_{j})$. The empirical parameter estimates align closely with the theoretical results.
@@
We demonstrated the bimodal posterior for Equation (2.1), and show that the behaviour can occur for different step sizes, priors, true parameters, numerical methods, and estimation methods. Where the second mode is located depends on the numerical method and step size, as which combination of bad parameters gives ‘correct’ $\hat{Y}(t_{j})$ depends on both those factors. Changes to the prior distribution were unable to appropriately constrain the second mode, but implementing the analytic solution instead of using a numerical method eliminated the erroneous parameters in MCMC sampling.
The persistence of bimodal posteriors due to numerical error indicate that choice of integration method matters beyond bias in estimating the posterior parameters. Bias has been tested in Agapiou et al., (2014) and O’Brien et al., (2024). We are able to select integration methods that produce unimodal estimates with negligible bias. For the affine model we used the analytic solution, and for the second non-linear DE known as the Canham function (Canham et al.,, 2004) RK45 with an adaptive step size produces accurate, unimodal posteriors.
[…]
Estimation methods are a consideration for models for both material (computation time) and theoretical reasons. We tested the computationally-intensive MCMC sampling method, and a computationally inexpensive but demonstrably less robust L-BFGS optimisation method that maximises the joint posterior likelihood. Neither avoided bimodality on their own because they too do not address the underlying numerical problems. MCMC sampling requires the numerical calculations to be run thousands of times so while an erroneous cluster can be identified after the fact with a high precision (and computationally expensive) numerical method, implementing that directly in the longitudinal model is not so feasible. That problems can be identified after the fact is helpful at least. In some cases multiple chains run in parallel can detect a bimodal posterior if they converge to different points, which is how we first observed the issue as tweaks to the von Bertalanffy model with RK4 failed to avoid the occasional chain converging to a secondary mode. Working with a small number of parallel chains is less reliable as a means of identifying posterior bimodality than large-scale simulation testing as we have done in this paper, particularly when the probability of converging to the second mode is small as for the affine model with step size 0.125.
@@
The numerical pathologies we found are adjacent to those addressed in the typical Bayesian workflow (Gelman et al.,, 2020), as the error comes from the longitudinal structure that is not part of most models. The simulation testing that we propose for posterior bimodality is relatively easy to implement, compared to hard-core numerical analysis for example, and can be done prior to large-scale dataset estimation. Aside from the affine model, we applied simulation tests to the extremely non-linear Equation (2.20), where we cannot apply an analytic solution to avoid bimodality as there isn’t one. In the case of the Canham function We are able to test for evidence of a bimodal posterior and parameter bias under the implemented numerical method – RK45 with adaptive step size – and find none. As a result we can be confident that the implementation of the Canham model using RK45 is unlikely to be as problematic as the von Bertalanffy model was without the analytic solution.

### v3

@@ 2.1 Numerical integration theory
We have described an example case where $\beta_{1}=1$, and $\beta_{0}=10$. The two real number solutions to Equation (8) (the complex solutions do not appear because parameter estimation is restricted to real values) for those values are $\hat{\beta}_{1}=1.0008$ and $4.9112$, which is shown in Figure 2 and aligns with the behaviour in Figure 1. The small difference from $\beta_{1}=1$ in the better estimate is to be expected given Corollary 2.1.
Figure 2: Plot of Equation (11), where the $y$-axis replaces 0 on the left hand side, with parameters $h=1/2$, $\beta_{1}=1$, and $\alpha=10$ for the first five steps. The intersections with 0 are at $\hat{\beta}_{1}=1.0008,\ 4.9112$.
To conclude the theory section, the posterior distribution for $\beta_{1}$ and $\beta_{0}$ conditioned on the RK4 step will have modes at the parameter values that produce $\hat{Y}(t_{j})$ that match the analytic solution. The solutions to Equation (8) gives posterior modes that cannot be distinguished by estimation based on a longitudinal model using the RK4 algorithm, which is precisely what non-identifiability means in this context.
Equation 8 is calculable directly because we have an analytic solution, explicit numerical method, and a straight-forward function. In other situations where either an analytic solution is unavailable, or if non-linearity makes an explicit step hostile to manipulation, simulation can provide empirical evidence of bimodality.
3 Empirical testing
In conjunction with the theoretical results that prove non-identifiability, we tested how the problem manifests in practice when estimating parameters from simulated repeat observation data.
@@ 3.1 Aims
Our key question is: Do different integration methods have persistent bimodality? We have demonstrated mathematical bimodality for RK4 with a step size of 0.5 so we test whether that shows up empirically as well. We also do empirical tests for RK4 with step sizes 0.25 and 0.125, RK45 with adaptive step size, and an analytic solution.
We have further questions are addressed in the supplementary material:
@@
  • Is there a model that provides accurate unimodal estimates? We do simulation studies for the linear function and a highly non-linear model to test whether the chosen work-arounds – an analytic implementation and RK45 with adaptive step size respectively – have reliable results.
@@ 3.2 Methods
To address the questions we simulated data so we know the true values. We produce ‘true’ data from the analytic solution Equation (3) with parameters $\beta_{0}=10$, $\beta_{1}=1$ and initial size $Y(0)=1$. The simulated survey structure is 10 size values $Y(t_{j})$ with an interval of 1 unit of time between them, so $t_{j+1}-t_{j}=1$. From the analytic sizes over time we get ‘observations’ with error, rounded to precision 0.1, of the form
@@
Figure 3(a) shows sizes over time with error compared to the analytic solution. The simulation structure was chosen to behave like observations in centimetres with precision of $1\text{\,}\mathrm{mm}$, and measurement error with $1\text{\,}\mathrm{mm}$ standard deviation. Results for the first two questions are based on running independent chains on the same set of 10 observations, so those results are conditioned on a single dataset. To test whether the implemented solution to bimodality is robust we simulate observations independently for each chain based on the same underlying $Y(t_{j})$.
@@
To avoid convergence problems we initialise the finite mixture algorithm with the sample means of the two initial clusters. The fitted mixture of normal distributions is then used to analyse the clustering behaviour, as we get cluster mean estimate vectors, covariance matrices, and cluster occurrence probability estimates. We also use scatter plots to look at the cluster marginal posterior parameter distributions. In the unimodal case we look at the mean of the estimated posterior parameters to check for evidence of large errors, and a 95% posterior credible interval (Gelman et al.,, 2021) built from taking the central 95% quantile interval of the posterior estimates to get an idea of spread.
@@
4 Results
4.1 Numerical methods have persistent bimodality
All the step sizes tested with the RK4 algorithm had a second cluster of erroneous parameter estimates. As the step size shrinks, the erroneous cluster moves further away from the true values, and the probability of erroneous convergence decreases as shown in Table 1. Figure 3 shows the cluster distance in Panel (b), where all 30,000 estimates are plotted together yet they appear as if single points due to the tiny within-cluster variance, as in Table 1. There is a stack of points at (9.95, 0.991) which is the first cluster of estimates for all step sizes, and each step size has a second cluster spread further from that point as the step size decreases. The clusters all lie along a line preserving the asymptotic size at $\hat{\beta}_{0}/\hat{\beta}_{1}\approx 10$, which is also visible in Figure 3(c) and (d). Panel (c) shows Equation (1) with parameters fit to the mean of each erroneous cluster compared to the true values, and Panel (d) shows the analytic solution Equation (3) with the same. The erroneous cluster has much faster analytic solution convergence to the asymptotic size, but as we saw in Figure 1, that leads to bad numerical behaviour.
[…]
We tested a better numerical method as well – RK45 with adaptive step size – and found that under some prior specifications bimodality still occurred. Figure 6 shows unimodal posteriors for both the analytic and RK45 posteriors when $\beta_{c}$ has a normal prior, while Figure 7 shows the bimodal posterior that arose when we was testing a log-normal distribution for $\beta_{c}$ with RK45, a problem that the analytic solution does not have. Log-normal priors are common so we include this example as a warning. Table 1 gives the mixture-model analysis for RK45. We do not consider the two clusters for $\beta_{c}\sim\log\mathcal{N}$ to be meaningfully different as they are so close together.
Table 1: Posterior finite mixture model estimates for MCMC testing different numerical methods for the linear ODE. True parameter values are $\beta_{0}=10$ and $\beta_{1}=1$. The sample standard deviations of each cluster for each parameter are given in parentheses. Note that for each choice of step size, RK4 converged to one of two very different solutions.
Numerical method Cluster 1 Cluster 2 Cluster 2
Step size $\mu_{\beta_{0},1}\ (sd(\hat{\beta}_{0},1))$ $\mu_{\beta_{1},1}\ (sd(\hat{\beta}_{1},1))$ $\mu_{\beta_{0},2}\ (sd(\hat{\beta}_{0},2))$ $\mu_{\beta_{1},2}\ (sd(\hat{\beta}_{1},2))$ prob.
@@ 4.1 Numerical methods have persistent bimodality
4.2 Implementing the analytic solution produces accurate, unimodal estimates
The simulation tests for the linear model with an analytic solution produced a unimodal distribution that was centred within 0.3% of the true values as shown in Table 2 which is a negligible empirical bias. Using the analytic solution means we do not introduce the non-identifiability that arises from numerical error so this result is in line with our theory.
Table 2: The linear model with an analytic solution shows minimal evidence of bias.
@@ 4.2 Implementing the analytic solution produces accurate, unimodal estimates
Linear $\beta_{0}$ 10 10.02 (9.484, 10.57)(9.484,\ 10.57)
$\beta_{1}$ 1 1.002 (0.9445, 1.062)(0.9445,\ 1.062)
@@ 5 Discussion
This paper provides theoretical and empirical evidence of a multi-modal posterior arising from numerical error in a longitudinal model for a linear first order ODE. One posterior mode aligns with good estimates of the true parameter values for simulated data, and at that mode the numerical method is reasonably well behaved. The other mode has extremely bad estimates of the true parameters, and badly behaved numerics as well. The existence of non-identifiability for RK4 in the case of a linear ODE demonstrates that care must be taken when using numerical methods for longitudinal models.
We show that the linear model can use an analytic solution to avoid the numerical pathology. The simulation testing procedure is more generally applicable and should form part of longitudinal model implementation if there is evidence of bad posterior behaviour such as divergent chains. It is also possible to test estimates from real world data after the fact by using a more accurate numerical method than the one implemented in the model.
Using numerical theory, we proved that in general, the RK4 algorithm in a longitudinal model for linear first-order ODEs with negative slope will produce multi-modal results, and demonstrate with a specific ODE. In the empirical results, we see two clusters of parameter estimates that produce statistically indistinguishable $\hat{Y}(t_{j})$. The empirical parameter estimates align closely with the theoretical results.
@@
We demonstrated the bimodal posterior for Equation (1), and show that the behaviour can occur for different step sizes and numerical methods. Where the second mode is located depends on the numerical method and step size, as which combination of bad parameters gives ‘correct’ $\hat{Y}(t_{j})$ depends on both those factors. Changes to the prior distribution were unable to appropriately constrain the second mode, but implementing the analytic solution instead of using a numerical method eliminated the erroneous parameters in MCMC sampling.
The persistence of bimodal posteriors due to numerical error indicate that choice of integration method matters beyond bias in estimating the posterior parameters. Bias has been tested in Agapiou et al., (2014) and O’Brien et al., (2024). We are able to select integration methods that produce unimodal estimates with negligible bias. For the linear model we used the analytic solution.
Our testing of different priors is distinct from the method recommended by Gabry et al., (2019). We seek to have default priors that resemble common choices to show that bimodality occurs, then test a specific set of priors to see if the problem can be averted. The choice of $\beta_{1}\sim\log\mathcal{N}(0,2)$ and $\beta_{c}\sim\mathcal{N}(1,2)$ would likely be ‘vague’ under the taxonomy proposed by Banner et al., (2020), though the standard deviation of 2 is much smaller than some choices for flat priors. Instead, the default standard deviation represents what could be plausible given the magnitude of measurement (on the order of 1 to $10\text{\,}\mathrm{cm}$) and the observed growth rates with a large-ish variance. The default priors are less important as we are not doing inference on data, W want to see if prior selection can constrain bimodality. We found that location did not, and the prior standard deviation needed to be unreasonably small to eliminate the second mode, far smaller than would ever be used in practice. We note that changing the prior distribution for the RK45 method from $\beta_{c}\sim\log\mathcal{N}(0,2)$ to $\mathcal{N}(1,2)$ did eliminate the second mode, so there may be specific cases where prior distribution choice works. The failure to eliminate the second mode with prior specification for the RK4 method is because the priors do not address the actual source of the problem.
Estimation methods are a consideration for models for both material (computation time) and theoretical reasons.MCMC sampling requires the numerical calculations to be run thousands of times so while an erroneous cluster can be identified after the fact with a high precision (and computationally expensive) numerical method, implementing that directly in the longitudinal model is not so feasible. That problems can be identified after the fact is helpful at least. In some cases multiple chains run in parallel can detect a bimodal posterior if they converge to different points, which is how we first observed the issue as tweaks to the von Bertalanffy model with RK4 failed to avoid the occasional chain converging to a secondary mode. Working with a small number of parallel chains is less reliable as a means of identifying posterior bimodality than large-scale simulation testing as we have done in this paper, particularly when the probability of converging to the second mode is small as for the linear model with step size 0.125.
@@
The numerical pathologies we found are adjacent to those addressed in the typical Bayesian workflow (Gelman et al.,, 2020), as the error comes from the longitudinal structure that is not part of most models. The simulation testing that we propose for posterior bimodality is relatively easy to implement, compared to hard-core numerical analysis for example, and can be done prior to large-scale dataset estimation.